OpenStudy (bloomlocke367):

An elliptical track has a major axis that is 80 yards long and a minor axis that is 72 yards long. Find an equation for the track if its center is (0,0) and the major axis is the x-axis.

OpenStudy (bloomlocke367):

so if the equation is \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] do I just square 80 and put it where a^2 is and square 72 and put it where b^2 is?

OpenStudy (bloomlocke367):

@iGreen @SolomonZelman

OpenStudy (krazykricket):

This question was already posted by someone on OS, here is the link to that question. http://openstudy.com/study#/updates/5166f830e4b0902dbc1db7f9

OpenStudy (phi):

a and b are semi-major and semi-minor i.e. 1/2 the length of the major axis and minor axis

OpenStudy (bloomlocke367):

I just figured that out. lol @phi

OpenStudy (bloomlocke367):

so is it \[\frac{ x^2 }{1600 } + \frac{ y^2 }{ 1296 } = 1\]

OpenStudy (phi):

yes

OpenStudy (bloomlocke367):

thanks :)